OpenStudy (anonymous):

Help Please I have two questions :) 1.)WXYZ is a parallelogram. Name an angle congruent to

OpenStudy (anonymous):



OpenStudy (timmyk14):

I believe the answer to your first question is D. angle WXY

OpenStudy (anonymous):

thank you :)

OpenStudy (anonymous):

2.) In parallelogram DEFG, DH=x+3, HF=3y, GH=4x-5, and HE=2y+3. Find the values of x and y. The diagram is not drawn to scale.

OpenStudy (anonymous):



OpenStudy (anonymous):

a.)x=6;y=3 b.)x=2;y=3 c.)x=3;y=2 d.)x=3;y=6

OpenStudy (timmyk14):

I forgot to explain why I chose answer choice D.. This is the definition of congruent: (of figures) identical in form; coinciding exactly when superimposed And angle WXY is identical to angle WZY

OpenStudy (anonymous):

I was thinking it was D too but I wasnt sure Thank You :)

OpenStudy (timmyk14):

For this you would simply plug in the values given (as the answers) and see which answer choice works out with this problem. So, which answer choice do you believe it is and I can show you whether or not it's right or not.

OpenStudy (anonymous):

oh this is too confusing....

OpenStudy (timmyk14):

Okay, then just guess and I'll help you work it out.

OpenStudy (anonymous):

okay um i think it may be a

OpenStudy (timmyk14):

Alright, so now we just have to plug in the values of x and y into the set of equations. Like this: DH = 6 + 3 HF = 3(3) GH = 4(6) - 5 HE = 2(3) + 3 Now we solve them... DH = 9 HF = 9 GH = 19 HE = 9 For me to correctly answer this question I need to know if the lines segments that comes to a point in the center of the figure are identical? By that I mean do they appear to be the same size in the picture you were given?

OpenStudy (anonymous):

yes the are identical, yes they are the same size

OpenStudy (timmyk14):

Okay, then answer choice A is wrong. Because they should all be the same value. Onto answer choice B.. Would you still like me to walk you through this or have you gotten the hang of it?

OpenStudy (anonymous):

could you maybe keep helping me i don't get geometry very well :)

OpenStudy (anonymous):

wait maybe i can get this

OpenStudy (timmyk14):

Okay, basically all we need to do know is solve the equation using the information given for answer choices B, C, and D. Wait, you don't need anymore help? Well, if you do just tag me by doing @Timmyk14

OpenStudy (anonymous):

Thank you for all your help :)

OpenStudy (timmyk14):

No problem :)